OpenStudy (explainitlikeimfive):

A manufacturer of shipping boxes has a box shaped like a cube. The side length is 5a + 4b. What is the volume of the box in terms of a and b? A. a^3 + 3a^2b + 3ab^2 + b^3 B. 125a^3 + 300a^2b + 240ab^2 + b^3 C. 125a^3 - 300 a^2b + 240ab^2 + 64b^3 D. a^3 - 3a^2b + 3a^2b - b^3

OpenStudy (anonymous):

Vol=l*w*h

OpenStudy (anonymous):

Vol of cube = s^3

OpenStudy (anonymous):

= (5a+4b)^3

OpenStudy (explainitlikeimfive):

I got B

OpenStudy (explainitlikeimfive):

Is that right?

OpenStudy (anonymous):

I got 125 a^3+300 a^2 b+240 a b^2+64 b^3

OpenStudy (anonymous):

V=side*side*side V=(5a+4b)(5a+4b)(5a+4b) V=(25a^2+40ab+16b^2)(5a+4b) V=125a^3+300a^2b + 240ab^2 + b^3 B.